Solve the following B.V.P.

𝑒𝑑𝑑 = 𝑐 2𝑒π‘₯π‘₯; { 𝑒(π‘₯, 0) = cos π‘₯ βˆ’ 1, 𝑒𝑑 (π‘₯, 0) = 0 𝑒(0,𝑑) = 0, 𝑒(2πœ‹,𝑑) = 0 0 ≀ π‘₯ ≀ 2οΏ½


1
Expert's answer
2021-11-01T11:48:57-0400

Given

"U_{tt}=2c\space U{xx}\\\frac{\delta ^2U}{\delta x^2}=\frac{1}{2c}\frac{\delta ^2 u}{\delta t^2}\\U(x,0)=f(x)=cos \space x-1\\U(x,0)=g(x)=0\\and\\U(0,t)=0,U(2\pi,t)=0\\0
≤ X ≤ 2"

We know that

"U(x,t)=\frac{1}{2}[f(x+ct)-f(x-ct)]+\frac{1}{2c}\int^{x+ct}_{x-ct}g(s)ds"

"U(x,t)=\frac{1}{2}[cos(x+\sqrt{2ct})-1-cos (x-\sqrt{2ct})+1]+\frac{1}{2\sqrt{2c}}\int ^{x+\sqrt{2ct}}_{x-\sqrt{2ct}}0\space ds"

"U(x,t)=\frac{1}{2}[cos(x+\sqrt{2ct})-cos (x-\sqrt{2ct})+1]"

and

"U(0,t)=0\\and\\U(2\pi,t)=0\\0=0"

we know that "cos\theta=cos \theta"

since verified

and "U(2\pi,t)=0"

"0=\frac{1}{2}[cos(2\pi+\sqrt{2ct})-cos(2\pi-\sqrt{2ct})]\\=\frac{1}{2}[cos(\sqrt{2ct})-cos(-\sqrt{2ct})]\\0=0"

verified

hence solution is

"U(x,t)=\frac{1}{2}[cos(x+\sqrt{2ct})-cos (x-\sqrt{2ct})]"